OpenStudy (anonymous):

Solve. 2x2 + 20x = –38 a. –5 ± √14 b. –5 ± √51 c. –5 ± √6 d. –5 ± √13

OpenStudy (anonymous):

c as usual. why is it always c?

OpenStudy (anonymous):

can u show me how, though, im trying hard to remember but its not coming to me...

OpenStudy (anonymous):

ok first off each term has a factor of 2, so we can make our life easier by writing \[x^2+10x=-19\]

OpenStudy (anonymous):

then we can write the left hand side as a perfect square , by taking half of 10, which is 5, squaring and adding it to both sides to get \[(x+5)^2=-19+25=6\]

OpenStudy (anonymous):

then take the square root, don't forget the plus and minus \[x+5=\pm\sqrt{6}\] then subtract 5 \[x=-5\pm\sqrt{6}\]

OpenStudy (anonymous):

oh i get it now, and this is for all problems like this?

OpenStudy (anonymous):

yes you can always use this method. if you like you could write \[x^2+10x+19=0\] and then use the quadratic formula, but that would be more work because you would have a denominator, and then you would have to simplify the radical, factor and cancel to get the answer you see above

OpenStudy (anonymous):

oh okay, thanks. can u also help me with http://openstudy.com/study?F295750376716BE3I5S=_#/updates/4f178119e4b0aeb795f57dc2

OpenStudy (anonymous):

if the coefficient of the "middle term" is even, then it is much easier to complete the square